OpenStudy (anonymous):

square root of 27 which decimal is approximately equal to: 2.56, 3.14, 4.2, 5.2

OpenStudy (anonymous):

you can go to http://en.wikipedia.org/wiki/Pi @mmehope

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

the square root of 27 is 5.19615242271, which rounds to 5.2

OpenStudy (anonymous):

can someone give me a medal plaese

OpenStudy (anonymous):

lol why

OpenStudy (anonymous):

you get a medal when you help someone with their problem

OpenStudy (anonymous):

i know i was jking with yall lol

OpenStudy (campbell_st):

well look at it this way in sequence you have 25, 27 and 36 25=5^2 and 36= 6^2 are square number... so the answer must be between 5 and 6

OpenStudy (anonymous):

@campbell_st is right